Circular No. 8081

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003az, 2003ba, 2003bb, 2003bc, 2003bd, 2003be Further to IAUC 8069, T. Dahlen, Space Telescope Science Institute,

on behalf of the same GOODS and Hubble Teams, reports the discovery of six new apparent supernovae on images of the Hubble Deep Field North, each detected by subtracting ACS images taken on 2002 Nov. 21–22 and 2003 Jan. 2–4 from those taken on 2003 Feb. 20–23, and each was independently detected in each of four separate z-band (F850LP) exposures, as well as in two separate i-band (F775W) exposures; none was detected in images from the earlier epochs. SN 2003ba, SN 2003bc, SN 2003bd, and SN 2003be were also detected in two separate v-band (F606W) exposures. The offsets given below are from the centers of the nearest visible galaxies. SN 2003bd has no apparent host galaxy to the detection limit of the survey (z ∼ 28). Further z-band magnitudes: SN 2003az, 2002 Nov. 22.13, [27; 2003 Jan. 3.85, [27. SN 2003ba, 2002 Nov. 22.40, [26; 2003 Jan. 4.84, [26. SN 2003bb, 2002 Nov. 21.93, [28; 2003 Jan. 2.30, [28. SN 2003bc, 2002 Nov. 21.33, [28; 2003 Jan. 2.51, [28. SN 2003bd, 2002 Nov. 21.76, [28; 2003 Jan. 2.91, [28. SN 2003be, 2002 Nov. 21.24, [27; 2003 Jan. 2.30, [27.

SN  2003 UT  α_{2000}  δ_{2000}  z  Offset
2003az Feb. 20.91 12^{h}37^{m}19.^{s}67 +62^{o}18′37′′.5 24.4 0′′.06 W, 0′′.08 S
2003ba Feb. 21.02 12 36 15.88 +62 12 37.7 23.7 0′′.18 W, 0′′.09 N
2003bb Feb. 21.63 12 36 24.47 +62 08 35.3 25.6 0′′.40 E, 1′′.31 S
2003bc Feb. 21.77 12 36 38.06 +62 09 53.3 24.0 1′′.10 W, 0′′.55 S
2003bd Feb. 21.95 12 37 25.06 +62 13 17.5 23.2 —
2003be Feb. 22.08 12 36 25.97 +62 06 55.6 23.1 0′′.12 W

X1901+031 Further to IAUC 8070, D. K. Galloway, R. Remillard, and E. Morgan

report an improved position for this recurrent x-ray transient, based on an RXTE/PCA scan on Feb. 12: α = 19^{h}03^{m}37^{s}.1, δ = +3^{o}11′31′′ (equinox 2000.0; estimated 1′ uncertainty at 90-percent confidence). This position is 1′.3 from the previous ASM position, and 9′ from the nearest ROSAT source, 1RXS J190411.9+031210. Measurements of the ∼ 2.763-s pulse period between Feb. 10 and 25 show variations suggesting orbital modulations with a period of ∼ 25 days or more. The ASM observations over the same time interval show a steady source with an average flux of 170 ± 5 mCrab and a hard spectrum with an ASM intensity (I) ratio of I(5–12 keV)/I(3–5 keV) = 2.6 ± 0.4. RXTE observations are continuing.

2003 February 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green